# DetailGen3D: Generative 3D Geometry Enhancement via Data-Dependent Flow Ken Deng1,2,3  Yuanchen Guo2  Jingxiang Sun1  Zixin Zou2  Yangguang Li2  Xin Cai4 Yanpei Cao2  Yebin Liu1  Ding Liang1,2

### v1

@@ 3.3 Token Matching
Due to the token representing the space area around the corresponding query points, we introduce a training method called token matching to match the input tokens with GT tokens as demonstrated in Fig 2. Without correct matching, the network not only learns the refine transformation but also location swapping because the tokens encoded by FPS-VAE are unordered.
Although attempts were made to match tokens in latent space, the correspondence is difficult to model effectively. In contrast, query points have explicit meaning. We explored the use of point cloud sorting algorithms, which work well for simple geometries but struggle with complex geometries due to their inability to preserve spatial relationships. Optimal Transport (OT) between coarse and fine query points can handle more complex shapes, but is computationally expensive and must be performed offline.
@@
To balance effectiveness and efficiency, we propose a novel token matching method. First, we obtain the query points of fine geometry by downsampling its point cloud. Next, we apply a nearest-neighbour algorithm to identify the closest fine geometry query points in the coarse geometry point cloud, using these as the coarse geometry’s query points. This method is computationally efficient and operates online, maintaining model generalization and offering robust performance across both simple and complex geometries, while ensuring local refinement.
Finally, it is important to note that replacing query points with learnable queries as proposed in [99] in the VAE does not address the issue. Firstly, the VAE’s performance with learnable queries is inferior to that when using query points. Secondly, learnable queries prevent the guarantee that tokens with the same indices in the encoded tensors of the coarse and fine geometries will correspond to similar features, leading to inconsistencies in token interpretation.
3.4 Dataset Design
To achieve local transformation between coarse and fine geometries, well-aligned coarse-fine pairs are essential. For coarse-fine pair design, as shown in Fig 2, we directly utilized large-scale reconstruction model reconstruction result of multi-view of fine model as coarse model. This approach minimizes the gap between prior observations and practical application. Unlike optimization-based methods, which are computationally expensive and time-consuming, this model completes reconstructions within 20 seconds, significantly improving dataset creation efficiency. The coarse-fine pairs produced by this method are well-aligned in both shape and spatial coordinates.
Although some geometric degradation, such as surface noise, remains for simple objects, our method has the advantage of inherently replicating the artifacts typical of neural network reconstructions. These artifacts are difficult to reproduce using human-designed priors. As a result, this method provides high-quality coarse-fine pairs that are both practically aligned and consistent with the models that will be encountered during inference, ensuring effective learning of the local transformation process.
@@ 4 Experiment
Our comprehensive evaluation of the model’s 3D geometric refinement capabilities is divided into three parts: 1). FeedForward-based Reconstruction: We tested the refinement ability of our method on the reconstruction results of various LRM models to demonstrate its effectiveness. 2). Generation: We assessed our method’s performance on the generation outputs from different LRM models, showcasing its versatility across diverse 3D representations. 3). Optimization-based Reconstruction: We applied our method to NeuS-like [75] optimization-based reconstruction, achieving impressive results that highlight its strong generalization. We also mixed results from different datasets, methods, and tasks for a user study, and conducted ablation studies on our design choices. See the supplementary for more details.
For the Objaverse [9] test set used, the IDs for models trained with different LRM methods are varied and not publicly available, which may lead to unfair comparisons between methods. However, we emphasize that our refinement model has never seen these 3D models during training, so its ability to refine the same object across different models fairly reflects the model’s generalization capability.
@@ 4.1 Feed-forward Reconstruction
Our task is to refine a given coarse geometry guided by images. To ensure that the coarse and refined geometries are spatially aligned to enable localized refinement, we use four GT views as input during training, obtaining a coarse model reconstructed by Instant3D [32]. Multi-view diffusion, while generating multiple views from a single image, can still introduce significant 3D inconsistencies, which severely affect reconstruction quality. In this experiment, we render the dataset according to the camera requirements of each LRM’s multi-view input, thus eliminating the impact of multi-view diffusion on geometry quality.
For the reconstruction task, we randomly sampled 199 models from the GSO [11] dataset and selected 350 unseen models from [9] as the evaluation dataset. We used Instant3D [32], CRM [80], InstantMesh [89] (testing both NeRF and Mesh, the checkpoints used in our work are all the official large versions provided.) as the reconstruction models for the coarse geometry. First, we rendered multi-view images of the GT according to the input viewpoint requirements of each large reconstruction model. This way, the reconstructed geometry can be assessed without the influence of different multi-view diffusion performances. We selected FID [21] as our evaluation metrics.
Visualization results can be found in Fig 3 and metrics results can be found in Tab 1. Our method fundamentally utilizes a generative model to add geometric detail to coarse geometry. Since only one view is used as guidance, this does not constrain the optimized geometry from being spatially consistent with the GT. Moreover, adding geometric detail without absolute spatial consistency with the GT can still increase loss. Therefore, we choose FID [21] as an evaluation metric. For the test images, we follow the setting of SDF-StyleGAN [105].
@@
Table 1: FID comparison of applying our method to different LRM results on reconstruction/generation task.
@@ 4.2 Generation
For the generation task, we randomly sampled 199 models from the GSO [11] dataset and selected 350 unseen models from Objaverse [9] as the evaluation dataset, same as mentioned before. Then used 110 images provided by GPTEval3D [86] as input, employing TripoSR [72], CRM [80] and InstantMesh [89] (testing both NeRF and Mesh, the checkpoints used in our work are all the official large versions provided.) as generation models for the coarse geometry. We selected FID [21] as our evaluation metrics. For the test images, we follow the setting of SDF-StyleGAN [105].
Although our method uses reconstruction results as input with GT as supervision, it can still generalize to non-reconstructed object, such as generated results. Our method shows outstanding performance in FID scores. As the visualization result illustrated in Fig 4 and metrics result shown in Tab 1, our method performs particularly well on complex objects, while previous methods lack robustness; with the same image input, they cannot guarantee consistently reasonable results across all generated models. Therefore, comparisons before and after refinement do not accurately reflect the capabilities of our model.
Compared to the reconstruction experiments, we conducted a more challenging test on the generation task. We used 110 images provided by GPTEval3D [86] as input. Due to the high level of difficulty, most single-view 3D generation models struggled to produce reasonable 3D models based on these images. We selected the plausible 3D models from this set and applied our refinement method. The results as illustrated in Fig 5 demonstrate the robustness of our approach to such challenging data.
4.3 Optimization-based Reconstruction
Figure 6: We apply our method on input meshes reconstructed by neus.   represent coarse,   represent fine refinement results from our method. The two objects on the left are from Objaverse, while the first object on the right is from GSO. More results can be found in supplementary.
@@ 4.3 Optimization-based Reconstruction
Table 2: FID scores evaluated on the Objaverse test set (350 objects) for different model ablations.
Figure 7: Visualization of ablation analysis on model designs.
In this experiment, due to the significant time cost of such methods, we were unable to test on the Objaverse [9] validation set (350 objects) and GSO [11] dataset (199 objects) as in the previous two experiments. Instead, we selected complex objects for reconstruction. For Neus [75], to further improve speed, we used Instant-NSR [17] implementation. For the multi-view data, we rendered a uniformly distributed set of 40 views as input.
Optimization-based reconstruction algorithms produce satisfactory details in multi-view reconstructions. However, due to the lack of prior guidance, they are prone to artifacts in scenarios with complex occlusions, and the smoothness of the surface remains imperfect. The refinement results as illustrated in Fig 6 demonstrate that our method successfully refines and corrects irregular geometric details.
@@ 4.4 User Study
Existing metrics do not fully capture the refinement level we aim to measure. We evaluated this by presenting four views (azimuths: 45°, 135°, 225°, 315°; elevation: 15°) of both coarse and refined geometries, rendered with consistent code. Images were shuffled, and volunteers voted on which geometry better matched the reference. Results show 94% of votes favored the geometry refined by our method.
@@ 4.5 Ablation Study
In our ablation study, we tested the effects of removing token matching, using cross attention to inject coarse geometry information, removing image prompts, removing noise augmentation, and using a point cloud sorting algorithm to query point cloud, inference without training for reference. We trained for approximately 100,000 steps on a dataset of 40,000 objects (subsample from the training set), with the results evaluated on a test set of 350 objects from Objaverse (same as in previous experiments). We used FID as the evaluation metric on the rendered images following the SDF-StyleGAN [105] settings. The FID results are shown in Tab 2, and the visualizations are presented in Fig 7.
Experimental results indicate that removing token matching slows convergence, as the model must simultaneously learn both refinement transformations and location swapping; it is hard to generate unseen details in the coarse geometry without the image condition; cross-attention modeling lacks the constraint of one-to-one token correspondence between coarse and fine models, making the noise-to-fine process inefficient and yielding worse geometric quality after refinement; applying point cloud sorting to the query point cloud yields no clear benefit, and inference without training is almost as effective in predicting coarse geometry.
Removing noise augmentation results in the lowest FID, as shown in Tab 2. We speculate that this is primarily due to two reasons: first, the Inception model [67], which computes FID, is not sensitive to the noise perception of the ensemble model, and second, the image conditioning ensures the results more consistent with the ground truth, even without noise augmentation. While these factors contribute to low FID, the visual degradation remains clearly noticeable, as demonstrated in Fig 7 and further illustrated in the supplementary material.
@@ 5 Discussion and Conclusion
Limitation and future work. As a generative model using a single image prompt, our approach cannot guarantee refinement in the exact direction of the ground truth. Future work will focus on enabling multi-view guided refinement to capture global information and improve alignment with the ground truth.

### v3

@@ 3.2 Token Matching
We propose token matching, a training method to match the latent codes of coarse and fine geometries for efficient training, as shown in Fig 2 (3). Training without token matching leads the network to learn not only refinement rules but also unnecessary operations (e.g., swapping latent codes), significantly reducing training efficiency. However, measuring the similarity between latent codes is non-trivial. Since the latent code represents the space area around the query points $X_{0}$ (Fig 2 (4)), we are inspired to match latent codes in latent space similarly to matching query points in 3D space.
@@
To balance effectiveness and efficiency, we propose a novel token matching method. First, we obtain the query points of fine geometry following 3D-VAE’s two-stage downsampling (applying random downsampling then farthest point sampling on fine geometry’s point cloud). Next, we obtain the coarse geometry’s query points by applying a nearest-neighbour algorithm. Specifically, for each fine query point $X_{0}^{\text{fine}}$, we identify its closest point within the full coarse point cloud. These identified points from the coarse cloud serve as the coarse query points $X_{0}^{\text{coarse}}$, establishing a spatial correspondence with $X_{0}^{\text{fine}}$. This method is computationally efficient, maintaining model generalization and offering robust performance across both simple and complex geometries while ensuring local refinement. We further discuss token matching in the supplementary material.
3.3 Data Curation
To achieve effective local transformation between coarse and fine geometries, well‐aligned coarse–fine pairs are crucial. Traditional geometric degradation techniques—such as Taubin smoothing [64]- can effectively reduce high‐frequency noise in complex geometries but tend to be too subtle for simpler ones. Moreover, applying the same degradation across all objects risks distorting complex shapes while failing to degrade simpler ones sufficiently. To address these issues, we adopt the reconstruction outputs from an LRM as shown in Fig 2 (2), which naturally introduce a balanced level of geometric degradation that adapts to an object’s inherent complexity. This choice not only preserves spatial correspondence but also better replicates the artifacts typical of neural reconstructions. Further details of data curation strategy are provided in the supplementary material.
Considering that fine objects are essential for training, we select Objaverse as our dataset and filter out the low-quality objects by evaluating each object’s quality based on CLIP [52] feature of four ortho rendering views. After filtering, our training set consists of about 110,000 objects, most of them are rich in details. We randomly partition the dataset into a training set and an evaluation set, with the latter containing 350 objects. The coarse geometry is reconstructed using a reimplemented version of Instant3D [31], leveraging four orthographic views of objects at a resolution of 512.
@@ 4 Experiment
Our comprehensive evaluation of the model’s 3D geometric refinement capabilities is divided into three parts: 1). FeedForward-based Reconstruction: We tested the refinement ability of our method on the reconstruction results of various LRM models to demonstrate its effectiveness. 2). Generation: We assessed our method’s performance on the generation outputs from different LRM models, showcasing its versatility across diverse 3D representations. 3). More 3D representations: We applied our method to reconstruction results reconstructed by NeuS-like [67] optimization-based and generation results generated by Rodin Gen-$1$https://hyperhuman.deemos.com/rodin and Neural4D [^2] https://www.neural4d.com/, achieving impressive results that highlight its strong generalization. We also performed ablation studies on our design choices. See the supplementary for more implementation and experiment details.
For the metric, we select FID [22] as our metric and obtain each object’s FID result following the setting of SDF-styleGAN [98], which can assess visual quality, capture statistical differences, detect mode collapse, and serve as a standardized benchmark despite not measuring fine-grained consistency perfectly. Considering FID may not fully capture the perceptible difference in fine details, we also mixed results from different datasets, methods, and tasks for a user study.
@@ 4.1 Feed-forward Reconstruction
Considering the coarse-fine pairs used in the training process, where all coarse models are reconstructed by Instant3D [31], this experiment aims to evaluate the generalizability to refining the reconstruction results of different kinds of LRM, including Instant3D as well.
For the evaluation set, we randomly sampled 199 models from the GSO [13] dataset and selected 350 unseen models from [11] as the evaluation set. For the LRM, We used Instant3D [31], CRM [71], InstantMesh [81] (testing both NeRF and Mesh, the checkpoints used are all the official large versions provided). We render multi-view images according to different LRM’s requirements and use their reconstruction results as coarse model.
Visualization results can be found in Fig 3 and FID results can be found in Tab 1. Our method can highlight the blurry details in the coarse model and add details only exist in the image prompt. In addition, for the aliasing existing on the input geometric surface, our method can eliminate them and show a smoother geometric surface. The FID measurement results reflect the effectiveness and robustness of our method for different models and different objects.
@@
Table 1: FID comparison of applying our method to different LRM results on reconstruction/generation task using Objaverse.
@@ 4.2 Generation
For the generation refinement task, we randomly sampled 199 models from the GSO [13] dataset and selected 350 unseen models from Objaverse [11] as the evaluation set, same as mentioned before. Additionally, to explore the refinement ability, we used 110 images provided by GPTEval3D [77] as input, employing TripoSR [65], CRM [71] and InstantMesh [81] (testing both NeRF and Mesh, the checkpoints used are all the official large versions provided.) as generation models for the coarse geometry.
As the visualization results illustrated in Fig 4 and FID results shown in Tab 1, our method performs particularly well on complex objects, while coarse models’ surface have large geometric noise and lack details. Furthermore, our method shows outstanding performance in FID scores.
For generation results, some of them are not well aligned with the input image. However, our method can still refine them in a proper way, where aligning the original shape and adding finer local detail. It shows that although our method learns refinement ability with aligned image-coarse shape pairs, our method knows how to use image prompt to add local details.
For experiments on GPTEval3D generation results, refinement results shows the robustness of our method to such challenging data, as illustrated in Fig 5. Each image from GPTEval3D [77] is used to obtain coarse models through image-to-3D models (i.e., TripoSR, CRM, InstantMesh) and serves as an image prompt in the refinement process. Due to the high level of difficulty, most image-to-3D generation models struggled to produce reasonable 3D models based on these images. We selected the plausible 3D models from them and applied our refinement method.
4.3 More 3D representations
Figure 6: We apply our method on input meshes reconstructed by NeuS and generated by Rodin Gen-1 and Neural4D.   represent coarse,   represent fine refinement results from our method. More results can be found in the supplementary.
@@ 4.3 More 3D representations
Table 2: FID scores evaluated on the Objaverse evaluation set (350 objects) for different model ablations.
Figure 7: Visualization of inference results with different settings.
To further evaluate the generalizability to other 3D representations, we tested our refinement ability on reconstruction results of NeuS [67] and generation results of two commercial products, i.e., Rodin Gen-1, Neural4D.
Considering the long optimization time of NeuS and the limited quota of Rodin Gen-1 and Neural4D, randomly sampled a subset of the evaluation set of Objaverse and GSO used in previous experiments for evaluation.
The refinement results are all illustrated in Fig 6 and demonstrate that our method successfully refines and corrects irregular details. The results of NeuS are reconstructed with 40 views evenly distributed around each object. However, due to the lack of prior guidance, they are prone to artifacts in scenarios with complex occlusions, and the smoothness of the surface remains imperfect. After refining by our method, the shredded parts become continuous, and details are more obvious. Since commercial products are typically trained on huge 3D data with a large number of GPUs, we selected images from GPTEval3D as input. The generation results have smooth surface but still lack details. After refinement, more details are introduced, and the shapes become more plausible.
@@ 4.4 User Study
The FID metric, while showing incremental improvement, may not fully capture the perceptible difference in fine details achieved by our method, especially on complex geometries. We invited 32 researchers, primarily experts in 3D vision, and compared the geometry quality of 10 models (reconstructed and generated by Instant3D [31], InstantMesh [81], CRM [71], TripoSR [65]) before and after refinement. In the questionnaire, we presented four views (azimuths: 45°, 135°, 225°, 315°; elevation: 15°) of both coarse and refined geometries, rendered with consistent code to ensure varied order and same color to avoid bias, and asked users to identify which of the two models (coarse vs. refined) appeared more detailed and closer to the input image. Statistics show that in 94% of the comparisons, participants judged the geometry obtained by our refinement method to be better (more detailed and closer to the image). This strongly validates the perceptual effectiveness of our method’s refinement ability.
@@ 4.5 Ablation Study
In our ablation study, we tested the effects of removing token matching, using cross-attention to inject coarse geometry information, removing image prompts, removing noise augmentation, and using a point cloud sorting algorithm to query point cloud, inference without training for reference. For all ablation experiments, we trained for approximately 100,000 steps on a dataset of 40,000 objects (subsample from the training set), with the results tested on an evaluation set of 350 objects from Objaverse (same as in previous experiments). We used FID as the evaluation metric on the rendered images following the SDF-StyleGAN [98] settings. The FID results are shown in Tab 2, and the visualizations are presented in Fig 7.
Experimental results indicate that removing token matching slows convergence, as the model must simultaneously learn both refinement transformations and unnecessary operation (e.g., position swapping); it is hard to generate unseen details in the coarse geometry without the image condition; cross-attention modeling lacks the constraint of one-to-one token correspondence between coarse and fine models, making the noise-to-fine process inefficient and yielding worse geometric quality after refinement; applying point cloud sorting to the query point cloud yields no clear benefit, and inference without training is almost as effective in predicting coarse geometry.
Removing noise augmentation results in the lowest FID, as shown in Tab 2. We speculate that this is primarily due to two reasons: first, the Inception model [61], which computes FID, is not sensitive to the noise perception of the ensemble model, and second, the image conditioning ensures the results are more consistent with the ground truth, even without noise augmentation. While these factors contribute to low FID, the visual degradation remains clearly noticeable, as demonstrated in Fig 7 and further illustrated in the supplementary material.
@@ 5 Discussion and Conclusion
[…]
Supplementary Material
Algorithm 1 Data-Dependent Rectified Flow
1:procedure $\mathcal{Z}$($\text{RectFlow}((X_{0},X_{1}))$)
2:       Inputs: Draws from a coupling $(X_{0},X_{1})$ of $\pi_{0}$ and $\pi_{1}$; velocity model $v_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with parameter $\theta$.
3:       Training:
4:       1. Obtain $\pi_{0}$ and $\pi_{1}$ using $\pi_{0}=\mathcal{E}(G_{0})$ and $\pi_{1}=\mathcal{E}(G_{1})$, where $\mathcal{E}$ denotes the VAE encoder, $G_{0}$ represents the coarse geometry, and $G_{1}$ represents the fine geometry.
5:       2. Optimize $\hat{\theta}=\arg\min_{\theta}\mathbb{E}\big{[}\|X_{1}-X_{0}-v(tX_{1}+(1-t)X_{0},t)\|^{2}\big{]}$, with $t\sim\text{Uniform}([0,1])$.
6:       Sampling:
7:       1. Start with $Z_{0}\sim\pi_{0}$, where $\pi_{0}=\mathcal{E}(G_{0})$. Here, $\mathcal{E}$ denotes the VAE encoder, $G_{0}$ represents the coarse geometry.
8:       2. Generate $(Z_{0},Z_{1})$ by solving $\mathrm{d}Z_{t}=v_{\hat{\theta}}(Z_{t},t)\mathrm{d}t$, obtaining $\{Z_{t}:t\in[0,1]\}$.
9:       Return: $\mathcal{Z}=\{Z_{t}:t\in[0,1]\}$.
10:end procedure
In this supplement, we first provide the implementation detail in Sec. 6. We also provide further experiment details in Sec. 7 and further discussion in Sec. 8. Finally, we provide additional visual results in Sec. 9. We encourage the readers to view our accompanying videos in the supplement, showcase the rotation of objects rendered with normals as presented in the paper.
6 Implementation Details
6.1 Model architecture
For the 3D-VAE encoder, after sampling $N$ points from geometry’s surface, we first randomly downsample it to $4\times M$, where $M$ is the number of latent codes of each object and $M=2048$, $N=20480$. Next, obtaining query points $X_{0}$ by using farthest point sampling to further subsample it with $\frac{1}{4}$ ratio.
The 3D-VAE decoder, comprised 24 multi-head self attention layer and a multi-head cross attention layer. For the preset query points used in cross attention layer, it evenly distributed in 3D space, using for querying the corresponding spatial SDF value, which can adopt marching cube algorithm to convert it into mesh format.
For the Refinement DiT is comprised 24 DiT blocks with a width of 768, 12 attention heads, and a latent length of 2048, totaling 368M parameters. For the image prompt, we use the image feature extracted by DINOv2 [47]. The feedforward network in each DiT block consists of a two layer multi layer perception with GELU activation and the middle dimension is four times the input dimension.
6.2 Noise Augmentation
During training, we apply noise to $z_{0}$ according to the DDPM [23] linear schedule at 400 timesteps. During the inference stage, noise augmentation is optionally, which can reduce the impact of floating objects in the coarse model on the final refined results. For our experiment, we all add noise at 100 time steps.
6.3 Data Curation
To construct satisfying coarse-fine pairs, we choose to obtain coarse model by using Instant3D to reconstruct through four ortho views at a resolution of 512. It is worth mention that some of the objects are textureless. For the training dataset, we select Instant3D [31] to obtain coarse model, and the objects are all from Objaverse [11]. It is worth noting that the Instant3D we use is reimplemented by us because its code has not been released yet.
In order to align the coarse geometry with the fine geometry in space, we first normalize and rescale each object to fit within a bounding box of side length 1, then translate the object so that the bounding box is centered at the origin.
6.4 Training
In our training setup, we start with a learning rate of 1e-10 and warm it up to 1e-4 over 5,000 steps. We use a total batch size of 256. We train our DetailGen3D model for 1,000 epochs, which takes approximately eight days on eight A800 GPUs. When training the data-dependent rectified flow, we randomly zero the DINOv2 features with a probability of 10% to enable classifier-free guidance during inference, thereby improving the quality of conditional generation. For the DINOv2 [47] checkpoint, we use the ViT-L/14 distilled with the registered version, downloaded from the official DINOv2 GitHub repository [^3] https://github.com/facebookresearch/dinov2. For the image prompt, we select a forward view with same camera intrinsic and extrinsic for training because using more views or highly random camera poses leads to longer training time. We believe that introducing camera pose embedding (e.g., plucker embedding) will help.
7 Experiment Details
Figure 8: Coarse model (left) is degraded and rotated manually following ShaDDR training setting. The model refined by ShaDDR (right) is worse.
7.1 Other Methods
We didn’t compare with other shape detailization methods as listed in related work because our tasks are different and they cannot handle our evaluation set, as illustrated in Fig.A 8. They aim at increasing the resolution of extremely low resolution voxels to relative higher resolution (e.g., $16^{3}$ to $64^{3}$). Our evaluation set consists of up to 450 objects, which are comes from GSO and Objaverse, however, their evaluation set only consists of tens objects and comes from the same category in ShapeNet [4] and have strict requirements on the orientation of objects (Refining a car from Objaverse using ShaDDR [6]’s checkpoint—which is trained on car objects from ShapeNet—produces worse results due to the difference in orientation compared to the training set.) and the input’s orientation shown in 8 is manually adjusted to meet ShaDDR’s requirements.
7.2 Evaluation Dataset
For the Objaverse [11] evaluation set used, the IDs for models trained with different LRM methods are varied and not publicly available, which may lead to unfair comparisons between methods. However, we emphasize that our refinement model has never seen these 3D models during training, so its ability to refine the same object across different models fairly reflects the model’s generalization capability.
7.3 SDF-stylegan Setting
For the feed-forward reconstruction and generation experiment, the FID [22] metric we evaluate following SDF-stylegan [98], rendering 20 views with preset random camera poses and same color (grey).
7.4 NeuS Reconstruction
For NeuS [67], to further improve speed, we used Instant-NSR [20] implementation. For the multi-view data, we rendered a uniformly distributed set of 40 views as input. The camera poses are elevation with -60°, -30°, 0°, 30°, 60° and azimuth with 0°, 45°, 90°, 135°, 180°, 225°, 270°, 315°.
8 Further Discussion
8.1 Token Matching
Although attempts were made to match latent code, the correspondence is difficult to model effectively. We explored the use of point cloud sorting algorithms, which work well for simple geometries but struggle with complex geometries due to their inability to preserve spatial relationships. Optimal Transport (OT) between coarse and fine query points can handle more complex shapes, but is computationally expensive and must be performed offline.
It is worth mentioning that using a learnable query in place of query points in the VAE does not solve the problem. The first reason is that the VAE quality with learnable queries is lower than that with query points. The second reason is that using learnable queries makes it impossible to ensure that the tokens with the same indices in the tensor, obtained after encoding the coarse geometry and fine geometry, have similar meanings.
Finally, we match the coarse geometry’s latent code with fine geometry’s latent code by applying nearest-neighbour algorithm to identify the closest fine geometry query points in the coarse geometry point cloud, using these as the coarse geometry’s query points. While this design cannot theoretically guarantee one-to-one correspondence, it does so experimentally. This is because the 3D-VAE’s two-stage downsampling process (i.e., downsampling point cloud $X$ sampled from geometry surface to query points $X_{0}$)—random downsampling followed by farthest point sampling—results in query points $X_{0}$ of fine geometry being distributed far apart in space. As a result, the nearest-neighbour algorithm becomes a viable method for matching latent codes.
8.2 Data Curation
While our experimental results have shown our method has strong generalizability across different sources of coarse models, including both generation and reconstruction tasks, there is still room for improvement. Using only one type of LRM (e.g., Instant3D) may introduce bias, and we believe that applying geometry degradation and mixing multiple LRM reconstruction results will be of help.
8.3 Application
Our focus on geometry refinement stems from the fact that in many applications (e.g., design, simulation) replying on fine geometry, whereas color can be integrated later [53, 75]. For the texture, our method supports to simply reproject the original textures onto the refined mesh or use separate texture-generation pipelines [90, 89]. This ensures high-quality geometry while preserving the flexibility to include color information as needed.
8.4 Limitation.
Our method can robustly refine geometry with image prompt. However, it still struggles with ultra-precision details like extremly complex and small geometry shown in image. This issue might be solved by training a better 3D-VAE, collecting more diverse coarse and fine geometry pairs for training.
9 Additional Visual Results
Due to the extensive nature of the content, the complete details and supplementary materials are available in the full appendix on our homepage. For further information, please visit:DetailGen3D